Value- and Non-Value-Added Cost Reporting

Cicleta Manufacturing has four activities: receiving materials, assembly, expediting products, and storing goods. Receiving and assembly are necessary activities; expediting and storing goods are unnecessary. The following data pertain to the four activities for the year ending 20x1 (actual price per unit of the activity driver is assumed to be equal to the standard price):

Activity Activity Driver SQ AQ SP
Receiving       Receiving orders 12,000     18,000     $21    
Assembly       Labor hours 75,000     90,000     15    
Expediting       Orders expedited 0     6,000     50    
Storing       Number of units 0     12,000     7    

Required:

1. Prepare a cost report for the year ending 20x1 that shows value-added costs, non-value-added costs, and total costs for each activity. If an amount is zero, enter "0".

Cicleta Manufacturing
Value and Non-Value-Added Cost Report
For the Year Ended 20X1
Activity Value-Added Costs Non-Value-Added Costs Total Costs
Receiving $ $ $
Assembly
Expediting
Storing
Total $ $ $
Answer #1
Value added costs non value added total
Receiving (12,000*21) = 252,000 (18,000-12,000)*21 = 126,000 378,000
Assembly (75,000*15) = 1,125,000 (90,000-75,000*15) = 225,000 1,350,000
Expediting 0 (6,000*50) = 300,000 300,000
Storing 0 (12,000*7) = 84,000 84,000
Total 1,377,000 735,000 2,112,000